# Tensor formalism for predicting synaptic connections with ensemble modeling or optimization

### v2

@@ VI.3.2 Characterizing the minimizing boundary using covariant and contravariant basis vectors
Figure 6: Geometric algorithm to determine $Q$-minimum and $Q$-critical in threshold linear network and related approximations. (A) We depict how the co- and contravariant basis vectors can be used to partition the semi-constrained subspace into the deep-yellow, orange, blue and yellow wedges spanned respectively by positive components of $(-\vec{\varepsilon}_{s1},-\vec{\varepsilon}_{s2})$, $(-\vec{\varepsilon}_{s1},\vec{\widetilde{\varepsilon}}^{s2})$, $(\vec{\widetilde{\varepsilon}}^{s1},\vec{\widetilde{\varepsilon}}^{s2})$, and $(\vec{\widetilde{\varepsilon}}^{s1},-\vec{\varepsilon}_{s2})$. (B) We illustrate the geometric algorithm to identify the correct boundary. We start from a point in the solution space, A, and move towards the location of the center, C. In the process we cross two wedge-boundaries, the first one flips $-\vec{\varepsilon}_{s2}\rightarrow\vec{\widetilde{\varepsilon}}^{s2}$, while the second $-\vec{\varepsilon}_{s1}\rightarrow\vec{\widetilde{\varepsilon}}^{s1}$, so that we know that C must be inside the blue wedge. (C-E) We simulate input-output transformations in feedforward networks of increasing size. (C) We plot the number of flips that the geometric algorithm required to reach the correct wedge. It seems to approximately grow linearly with the size of the network. (D) We depict the norm of the optimized weight vectors obtained using quadratic programming (light blue) and the geometric algorithm (dashed black), as the network size increases. (E) We plot the small angular difference between the two optimized vectors. (F) We plot the $Q_{\mathtt{\min}}$ values corresponding to the optimized weight vectors obtained using quadratic programming (light blue) and the approximate weight vectors (blue and dashed red) obtained using natural, Eq. (102), and check complementary approximation Eq. (103), schemes respectively, as the network size increases. The grey region depicts the range of values $Q_{\mathtt{\min}}$ must be restricted to according to the analytically obtained bound given by Eq. (108). (G) The difference in the corresponding $Q$-minimum values between the optimized (obtained numerically using quadratic programming) and approximate vectors are plotted. To facilitate comparison between the two different approximations we plot the difference in $Q$ values coming from the two approximate vectors against one another. (H) Same as in (G), except that we compare the Euclidean angular error corresponding to the natural and complementary approximation.
@@ VI.3.4 An approximation to determine $Q_{\mathtt{\min}}$
then $\vec{c}$ doesn’t belong to the solution space, as the network specified by $\vec{c}$ will then evoke a positive response, $\vec{c}\cdot\vec{\varepsilon}^{\,\widetilde{\mu}}>0$, to the $\mu^{th}$ pattern. One has to therefore consider weight vectors in the semi-constrained subspace that have norms larger than $||\vec{c}||$. As argued previously, the most judicious choice is to keep increasing the norm till we find a vector, $\vec{s}$, whose projection is zero along $\vec{\varepsilon}^{\,\widetilde{\mu}}$, effectively turning $\widetilde{\mu}$ into a constrained dimension so the the network just starts to not respond to the pattern. We know that this argument breaks down when the patterns are non-orthogonal but we surmise that finding the $W_{\mathtt{\min}}$ for a linear problem where all the semi-constrained dimensions with $\vec{c}\cdot\vec{\varepsilon}^{\,\widetilde{\mu}}>0$ are treated as constrained along with the constrained patterns may provide a good approximation to the real $W_{\mathtt{\min}}$ and $\vec{w}_{\min}$. Further, we decided to extend our hypothesis to the general problem of finding $Q_{\mathtt{\min}}$ by covariantly generalizing the conjecture (102) to inner products given by $g_{mn}=Q_{mn}$.
@@
where $\gamma_{\dot{\mu}\dot{\nu}}$, $\gamma_{{\cal B}_{0},\dot{\mu}\dot{\nu}}$, and $\gamma_{\emptyset,\dot{\mu}\dot{\nu}}$ are the metric components corresponding to the full weight-space, the induced subspace containing all the constrained dimensions and the semi-constrained dimensions used by the natural or complementary approximations (see text below Eq. (104), and the constrained subspace, respectively. We’ll soon see (Fig. 6F-H) that these approximations and bounds can be quite accurate.
@@ VI.3.5 Numerical verification
In Fig. 6C we have plotted the number of flips that was required to arrive at the solution as the size of the network increased. As we suspected, the growth seems linear strongly suggesting that our algorithm doesn’t suffer from a combinatorial explosion. To verify the accuracy of our geometric approach we decided to compare our results with solutions that can be obtained using standard convex optimization approaches to quadratic programming Boyd and Vandenberghe (2004); Wright (1999). In Fig. 6D,E, we compare our results using the two different methods. In Fig. 6D we have plotted the norm of the optimum weight vectors, the norms from geometric and quadratic programming are plotted in light blue and dashed black respectively, and seem indistinguishable. Fig. 6E shows the angular difference between the two vectors, it is ${\cal O}(10^{-2})$ degrees, further confirming the agreement between the two methods. Our geometric approach thus might provide a new and insightful way to solve quadratic programming problems, which have wide applicability McCarl et al. (1977).
In the previous subsection we also obtained $\vec{w}_{\min}$, Eq. (104), and $Q_{\mathtt{\min}}$ Eq. (108) using natural and complementary approximation schemes. In Fig. 6F-H, we try to assess how well the approximation schemes work. In  6F, we plot the $Q_{\mathtt{\min}}$ values obtained using quadratic programming and the two approximation schemes along with the $Q_{\mathtt{\min}}$ range given by the upper and lower bounds. In Fig. 6H we compute the error or difference in the $Q_{\mathtt{\min}}$ values obtained using quadratic programming and the two approximations. The average discrepancy is about a few percents. Fig. 6H is similar, except that we compute the Euclidean angles, $\theta_{\mathtt{diff}}$’s, between the optimized vector and the approximate vectors according to,

### v3

@@ VI.3.2 Characterizing the minimizing boundary using covariant and contravariant basis vectors
Figure 6: Geometric algorithm to determine $Q$-minimum and $Q$-critical in threshold linear network and related approximations. (A) We depict how the co- and contravariant basis vectors can be used to partition the semi-constrained subspace into the deep-yellow, orange, blue and yellow wedges spanned respectively by positive components of $(-\vec{\varepsilon}_{s1},-\vec{\varepsilon}_{s2})$, $(-\vec{\varepsilon}_{s1},\vec{\widetilde{\varepsilon}}^{s2})$, $(\vec{\widetilde{\varepsilon}}^{s1},\vec{\widetilde{\varepsilon}}^{s2})$, and $(\vec{\widetilde{\varepsilon}}^{s1},-\vec{\varepsilon}_{s2})$. (B) We illustrate the geometric algorithm to identify the correct boundary. We start from a point in the solution space, A, and move towards the location of the center, C. In the process we cross two wedge-boundaries, the first one flips $-\vec{\varepsilon}_{s2}\rightarrow\vec{\widetilde{\varepsilon}}^{s2}$, while the second $-\vec{\varepsilon}_{s1}\rightarrow\vec{\widetilde{\varepsilon}}^{s1}$, so that we know that C must be inside the blue wedge. (C-E) We simulate input-output transformations in feedforward networks of increasing size. (C) We plot the number of flips that the geometric algorithm required to reach the correct wedge. It seems to approximately grow linearly with the size of the network. (D) We depict the norm of the optimized weight vectors obtained using the quadprog function in Matlab (light blue) and the geometric algorithm (dashed black), as the network size increases. (E) We plot the small angular difference between the two optimized vectors. (F) We plot the $Q_{\mathtt{\min}}$ values corresponding to the optimized weight vectors obtained using quadprog function (light blue) and the approximate weight vectors (blue and dashed red) obtained using natural, Eq. (102), and check complementary approximation Eq. (103), schemes respectively, as the network size increases. The grey region depicts the range of values $Q_{\mathtt{\min}}$ must be restricted to according to the analytically obtained bound given by Eq. (108). (G) The difference in the corresponding $Q$-minimum values between the optimized (obtained numerically using quadprog function) and approximate vectors are plotted. To facilitate comparison between the two different approximations we plot the difference in $Q$ values coming from the two approximate vectors against one another. (H) Same as in (G), except that we compare the Euclidean angular error corresponding to the natural and complementary approximation.
@@ VI.3.3 Geometric prescription to address the combinatorial challenge in identifying the boundary
The fact that the tensor formalism that was developed to describe curve space-time in physics could, in fact, be used to provide insights into a biological problem was somewhat surprising to us. We believe that the geometric tensor-based framework we have developed will enable us to go beyond quadratic programming and, for instance in a manner similar to what was discussed in the context of orthogonal patterns Biswas and Fitzgerald (2022), help us deal with noise and modifications of input-output transformation functions that are not threshold linear. Whether our geometric formalism based on distance minimization can be generalized to deal with optimizing functions more general than quadratic, remains an open question, but this was also an important motivation for proving a link between physics-based tensor formalism and biologically relevant optimization/ensemble-based predictions.
@@ VI.3.4 An approximation to determine $Q_{\mathtt{\min}}$
then $\vec{c}$ doesn’t belong to the solution space, as the network specified by $\vec{c}$ will then evoke a positive response, $\vec{c}\cdot\vec{\varepsilon}^{\,\widetilde{\mu}}>0$, to the $\mu^{th}$ pattern. One has to therefore consider weight vectors in the semi-constrained subspace that have norms larger than $||\vec{c}||$. As argued previously, the most judicious choice is to keep increasing the norm until we find a vector, $\vec{s}$, whose projection is zero along $\vec{\varepsilon}^{\,\widetilde{\mu}}$, effectively turning $\widetilde{\mu}$ into a constrained dimension so the the network just starts to not respond to the pattern. We know that this argument breaks down when the patterns are non-orthogonal but we surmise that finding the $W_{\mathtt{\min}}$ for a linear problem where all the semi-constrained dimensions with $\vec{c}\cdot\vec{\varepsilon}^{\,\widetilde{\mu}}>0$ are treated as constrained along with the constrained patterns may provide a good approximation to the real $W_{\mathtt{\min}}$ and $\vec{w}_{\min}$. Further, we decided to extend our hypothesis to the general problem of finding $Q_{\mathtt{\min}}$ by covariantly generalizing the conjecture (102) to inner products given by $g_{mn}=q_{mn}$.
@@
where $\gamma_{\dot{\mu}\dot{\nu}}$, $\gamma_{{\cal B}_{0},\dot{\mu}\dot{\nu}}$, and $\gamma_{\emptyset,\dot{\mu}\dot{\nu}}$ are the metric components corresponding to the full weight-space, the induced subspace containing all the constrained dimensions and the semi-constrained dimensions used by the natural or complementary approximations (see text below Eq. (104), and the constrained subspace, respectively. We’ll see (Fig. 6F-H) that these approximations and bounds can be quite accurate.
@@ VI.3.5 Numerical verification
In Fig. 6C we have plotted the number of flips that was required to arrive at the solution as the size of the network increased. As we suspected, the growth seems linear, strongly suggesting that our algorithm doesn’t suffer from a combinatorial explosion. To verify the accuracy of our geometric approach, we decided to compare our results with solutions that can be obtained using standard convex optimization approaches to quadratic programming Boyd and Vandenberghe (2004); Wright (1999). In Fig. 6D,E, we compare our results using the two different methods. In Fig. 6D we have plotted the norm of the optimum weight vectors, the norms from geometric and quadratic programming are plotted in light blue and dashed black respectively, and seem indistinguishable. Fig. 6E shows the angular difference between the two vectors, it is ${\cal O}(10^{-2})$ degrees, further confirming the agreement between the two methods. Our geometric approach thus might provide a new and insightful way to solve quadratic programming problems, which have wide applicability McCarl et al. (1977).
In the previous subsection we also obtained $\vec{w}_{\min}$, Eq. (104), and $Q_{\mathtt{\min}}$ Eq. (108) using natural and complementary approximation schemes. In Fig. 6F-H, we try to assess how well the approximation schemes work. In  6F, we plot the $Q_{\mathtt{\min}}$ values obtained using quadratic programming and the two approximation schemes along with the $Q_{\mathtt{\min}}$ range given by the upper and lower bounds. In Fig. 6H we compute the error or difference in the $Q_{\mathtt{\min}}$ values obtained using quadratic programming and the two approximations. The average discrepancy is about a few percent. Fig. 6H is similar, except that we compute the Euclidean angles, $\theta_{\mathtt{diff}}$’s, between the optimized vector and the approximate vectors according to,
@@
In Fig. 8 in Appendix F, we also compare approximate run times of a standard quadratic programming implementation (the quadprog function with default parameters in MATLAB) to our geometric algorithm and analytical approximations (also in MATLAB). We found that our analytical approximation was fastest for all tested problems. The geometric algorithm’s run times were faster than quadprog up to a network size of ${\cal N}\sim{\cal O}(100)$, but quadprog ran faster than the geometric algorithm as the network size increased.